OpenStudy (anonymous):

Is anyone good at Point-Slope form, and stuff like that? With perpendicular lines?

OpenStudy (solomonzelman):

\[\huge {Shoot!}\]

OpenStudy (anonymous):

Ok, can you help me solve this? Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3

OpenStudy (solomonzelman):

You have y = 2x + 6 the perpendicular slope for 'm' is going to be -1/m, this is always true. So for 2, it is -1/2. now rewrite it, y = -1/2x +b we know the slope of the perpendicular equation, but we have to find the y-intercept plug your point in (6, 3) y = -1/2x +b ---> (3) = -1/2(6) +b ----> 3=-3+b add 3 to both sides. 6=b so you have y = -1/2x + 6 as your final answer.

OpenStudy (solomonzelman):

@Niall, \[\huge {Good?}\]

OpenStudy (anonymous):

Yeah I think I get it. I have another more confusing one, can you help me with it too?

OpenStudy (solomonzelman):

I can try, I am not good at math.

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

OpenStudy (solomonzelman):

3x + y = 7 ---> into slope intercept form ---> y = -3x+7 You have y = -3x + 7 the perpendicular slope for 'm' is going to be -1/m, this is always true. So for -3, it is 1/3. now rewrite it, y = 1/3x +b we know the slope of the perpendicular equation, but we have to find the y-intercept plug your point in (6, −1) y = 1/3x +b ---> (-1) = 1/3(6) +b ----> -1=2+b subtract 2 from both sides. -3=b so you have y = 1/3x - 3 as your final answer.

OpenStudy (solomonzelman):

See it is the SAME THING!

OpenStudy (anonymous):

Ohh ok. I'll post another 1, but don't give me the answer. I'm gonna try to solve it first. Then tell me if its right please. :)

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = −three fourthsx + 8 y = −three fourthsx + 3

OpenStudy (anonymous):

Ok, is it y = 4/3x + 8?

OpenStudy (solomonzelman):

can't be right, it is wrong.

OpenStudy (anonymous):

Dang, do you know what it is?

OpenStudy (solomonzelman):

yes, I can show you, you got the right slope, but not the y intercept. y = 4/3x + b plug in (−3, 4) into this. And you were right! SORRY!

OpenStudy (solomonzelman):

I miscalculated, you were correct!

OpenStudy (anonymous):

Yayyyy! :P Ok, thanks a lot for the help! :)

OpenStudy (solomonzelman):

\(\huge {ANYTIME!}\) \(\huge {(•‿•)}\)